# NEAT: Neighborhood-Guided, Efficient, Autoregressive Set Transformer for 3D Molecular Generation

### v1

@@ Abstract
Autoregressive models are a promising alternative to diffusion-based models for 3D molecular structure generation. However, a key limitation is the assumption of a token order: while text has a natural sequential order, the next token prediction given a molecular graph prefix should be invariant to atom permutations. Previous works sidestepped this mismatch by using canonical orders or focus atoms. We argue that this is unnecessary. We introduce NEAT, a Neighborhood-guided, Efficient, Autoregressive, Set Transformer that treats molecular graphs as sets of atoms and learns the order-agnostic distribution over admissible tokens at the graph boundary with an autoregressive flow model. NEAT approaches state-of-the-art performance in 3D molecular generation with high computational efficiency and atom-level permutation invariance, establishing a practical foundation for scalable molecular design.
Machine Learning, ICML
@@ 1 Introduction
Computational workflows for discovering bioactive molecules often rely on either virtual screening of compound libraries or fragment spaces (warr2022exploration), or direct de novo generation of compounds with desired properties (Peng2023MolDiff). Although the expansion of chemical spaces increases the odds that the right design will be found, their growing sizes make screening increasingly computationally expensive (bellmann2022comparison). Moreover, despite their ultra-large sizes, chemical spaces still encompass only a small fraction of the vast molecular design space.
@@
Conditional molecular generation offers a promising alternative to virtual screening by enabling the computationally cheaper generation of new molecules based on user-defined constraints. It can be approached through the generation of (1D) SMILES strings (GomezBombarelli2018CVAE; Kusner2017GVAE), (2D) molecular graphs (Simonovsky2018GraphVAE; DeCao2022MolGAN), (3D) atomic point clouds (Gebauer2019GSchnNet; GarciaSatorras2021ENF; Hoogeboom2022EDM; Luo2022GSphereNet; Xu2023GeoLDM; Daigavane2024Symphony; Zhang2025SymDiff; Cheng2025Quetzal), or a combination thereof (Hua2023MuDiff; Peng2023MolDiff; Vignac2023MiDi). Because molecules are inherently three-dimensional, with many properties determined by their conformations, 3D generation is the self-evident choice. In this work, we focus on unconditional 3D generation, which is a necessary prerequisite to condition the generation process towards compounds with specific biological functions later on (peng2022pocket2mol; schneuing2024structure).
@@
State-of-the-art 3D molecular generation models rely on diffusion and flow matching techniques (Hoogeboom2022EDM; Xu2023GeoLDM; Song2024GeoBFN; Zhang2025SymDiff; Feng2025UniGEM; Song2023EquiFM; Dunn2025FlowMol). A key limitation of these methods is the need to compute the joint atom distribution at each diffusion step through computationally intensive, quadratically scaling message-passing updates in the encoder model to capture atom interdependencies. Additionally, diffusion and flow matching molecular generators require the number of atoms to be predefined. While the use of dummy atoms offers a partial workaround (adams2025shepherd; Dunn2025FlowMol), these approaches remain constrained by the upper bound on the number of nodes sampled from the noise distribution.
@@
Autoregressive models offer a flexible and efficient alternative. They support variable atom counts and require only one encoder call per generated token (Li2024AutoregressiveImageGenerationWithoutVQ). For 3D autoregressive molecular generation, there exists 3 lines of work: (1) using graph neural networks (GNNs) in combination with focus atoms (Gebauer2019GSchnNet; Luo2022GSphereNet; Daigavane2024Symphony); (2) using standard large language models (LLMs) for generating XYZ, CIF or PDB files (flamshepherd2023llm2mol) or SE(3)-invariant 1D tokens (Li2025Geo2Seq; Fu2025Frag2Seq; lu2025octree); (3) combining a causal transformer with a diffusion loss to generate continuous tokens (atomic coordinates) (Cheng2025Quetzal; Li2025InertialAR). Although current GNN-based methods ensure permutation invariance with respect to atom ordering, they generate molecules with low chemical validity. Conversely, sequence-based methods perform well in terms of validity but are not permutation-invariant with respect to atom ordering.
@@
In this work, we present the first autoregressive transformer model for 3D molecular generation that respects the atom permutation-invariance of molecular graphs. Working without canonical atom orderings introduces two challenges. First, it precludes the use of sequence embeddings, which are an essential component of standard autoregressive models. To address this, we model atomic interdependencies within the growing molecular graph using a set transformer (Lee2019SetTransformer). Second, it changes the autoregressive training set-up from supervision with a single token per step to supervision with a set of valid tokens: without a canonical atom order, all atoms in the 1-hop neighborhood of the current graph are valid continuations.
@@
Our solution to this challenge is conceptually simple and effective. Given a partial molecular graph, the model is trained to produce a token distribution that matches the boundary nodes of the subgraph (i.e., the 1-hop neighbors of existing nodes). We refer to this training signal as neighborhood guidance. This yields a probabilistic next-token distribution without relying on ad hoc canonicalization schemes. We combine this with a flow-matching loss to model continuous atomic coordinates. Our model, NEAT (Neighborhood-guided, Efficient Autoregressive set Transformer), achieves competitive performance relative to state-of-the-art 3D generative models while maintaining permutation invariance to atom ordering.
We expect these contributions to advance the development of autoregressive methods for de novo 3D molecular design and to provide a solid foundations for conditional molecular generation.
@@ 2 Related works
$p(x)=\prod_{n=1}^{N-1}p(x_{n+1}\mid x_{:n}),$ (1)
where $x_{:n}=(x_{1},\dots,x_{n})$ represents the tokens up to step $n$. Sampling is performed by initializing with $x_{1}$ and sequentially generating tokens until $x_{N}$. This is straightforward for sequentially ordered data such as text or images. In the graph domain, the generation order is less obvious, and has been an active focus of research. One way to model graphs as sequences is through canonicalization. However, predefined canonical orderings impose limitations, as the optimal generation order is often context-dependent. To alleviate strict assumptions, you2018graphrnn proposed a breadth-first search (BFS) node ordering for autoregressive graph generation, while wang2025learningorder introduced order policies to dynamically assign generation orders based on the current state. Although these approaches reduce reliance on fixed orderings, they do not fully eliminate order dependence. 3D molecules are geometric graphs whose structure is determined by connectivity and geometry. The generation of valid positions poses additional constraints on the graph generation process and recent work on this specific task will be discussed in the following.
@@
G-SchNet (Gebauer2019GSchnNet) is an $E(3)$-invariant GNN model that introduces the concept of focus atoms for generation order. In each generation step, the model selects a focus atom as the center of a 3D grid, chooses the next atom type, and infers its position probability from radial distances to previously placed atoms. G-SphereNet (Luo2022GSphereNet) adopts this approach but places atoms relative to the focus atom by generating distances, angles, and torsions with a normalizing flow. The $E(3)$-equivariant Symphony model (Daigavane2024Symphony) parameterizes the 3D probability around the focus atom using $E(3)$-equivariant features from spherical harmonic projections rather than rotationally invariant features. While these models respect the permutation-invariance of atoms, they produce molecules with comparably low chemical validity.
Transformer models with tokenization
Transformers can be used for 3D molecule generation when combined with a suitable text format. flamshepherd2023llm2mol trained an LLM for 3D molecular generation directly on XYZ, CIF, or PDB files. An active line of research are tokenization schemata to allow LLMs more fine-grained control in the generation process. Geo2Seq (Li2025Geo2Seq) converts molecular geometries into $SE(3)$-invariant 1D discrete sequences for unconditional modeling, while Frag2Seq (Fu2025Frag2Seq) introduces a related technique for structure-based modeling. lu2025octree employs a voxel-based discretization scheme with an octree-based compression technique, yielding sparse sequences.
@@
One of the challenges in generating 3D molecular structures with transformers is to generate continuous-valued tokens like 3D atomic coordinates without discrete tokenization strategies. While standard transformer architectures were initially developed for categorical data, they can be adapted for the generation of continuous values. Li2024AutoregressiveImageGenerationWithoutVQ directly models the next-token distribution $p(x_{n+1}\mid x_{:n})$ using a diffusion-based approach, factorized as:
$p(x_{n+1}\mid x_{:n})=p(x_{n+1}\mid\mathbf{z}_{n})\cdot p(\mathbf{z}_{n}\mid x_{:n}),$ (2)
where $\mathbf{z}_{n}\in\mathbb{R}^{k}$ represents an abstract latent variable for the next token. This factorization enables to use a transformer to model $p(\mathbf{z}_{n}\mid x_{:n})$ to capture interdependencies among tokens, followed by a lightweight diffusion-based MLP to map $\mathbf{z}_{n}$ to the continuous token space. Their approach demonstrated high-quality autoregressive image generation with improved inference speed.
Recent work successfully applied the diffusion loss to the generation of atomic positions, enabling 3D molecular generation without tokenization. Cheng2025Quetzal employed a causal transformer with sequential embeddings derived from the atom order in XYZ coordinate files to predict discrete atom types, and a diffusion model, conditioned on these predictions, to predict the corresponding continuous 3D atom positions. However, the study also shows that reliance on a causal transformer introduces limitations. The use of sequence information based on atom order in XYZ files makes the model susceptible to overfitting and degrades performance when the atom order is perturbed. (Li2025InertialAR) adopt a comparable framework, adding inertial-frame alignment for coordinates and canonical atom ordering, but their method only relocates the problem of token-order sensitivity rather than eliminating it.
We propose that a set transformer (Lee2019SetTransformer) with bidirectional attention is better suited for autoregressive 3D molecular generation, as it inherently respects the permutation invariance of atom sets by disregarding generation order at state $n$. This raises the challenge of compensating for the absence of sequential order information, which is the key focus of our work.
@@ 3 Methods
Our architecture consists of three main components: (1) a data-augmentation module that provides neighborhood-guidance during training, (2) a set transformer trunk that produces permutation-invariant encodings of atom sets, and (3) a flow-matching head that generates continuous atomic coordinates. During training, we split molecules into a source set, representing the state of the molecular graph at step $n$, and a target set, representing the set of possible token continuations at step $n+1$. The source set is processed by the set transformer to obtain a learned representation of the current state. This source set representation is then passed through a simple linear linear layer followed by a softmax to yield a categorical distribution over the next atoms’ types. The continuous next atoms’ positions are modeled by the flow matching head, conditioned on the source set representation and the next atom type. During training, we use teacher forcing to decouple the prediction of atom types from the prediction of atom positions. Because the expensive source set representation is computed only once per generated token, our model is computationally more efficient than non-autoregressive frameworks that rely entirely on diffusion or flow-matching. Training and inference procedures are summarized in Figures 1 and 2, respectively.
@@
Figure 1: Overview of the proposed training workflow. The model takes molecular graphs as input, where nodes encode atom types and positions, and edges represent chemical bonds. For each training example, we randomly sample a connected subgraph. The selected nodes form the source set, and their neighboring boundary nodes define the target set. The source set is encoded using a set transformer, and its representation is used to predict the distribution over the next atom type. For each target-set node, we sample a position from a normal distribution and construct linear interpolation paths between the sampled and true positions. These interpolation paths provide supervision for learning the vector field via flow matching.
@@
In a build-up phase, the model iteratively predicts atom type $a_{n+1}$ and position $\mathbf{x}_{n+1}$, given the current set of atoms and positions at step $n$. We will use the following compact notation $\mathbf{a}_{1:n}=(a_{1},\dots,a_{n})$ to refer to the generated atom types and positions $\mathbf{X}_{1:n}=(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ at step $n$. The autoregressive build-up stops when the model predicts a stop token. Edges are then inferred from the $N$ generated atoms via a predefined set of rules like $E=\mathcal{R}(\mathbf{a},\mathbf{X})$, which completes inference of the molecular graph $G$. While edges are inferred from atom types and positions after the build-up phase, the edge information can still be used for teacher guidance in the training phase. In the following, we describe how we use this information for model training.
@@
We propose a sampling strategy that mimics molecular growth during the autoregressive build-up phase. Given a molecule $G$ in the training set, we sample a node-induced, connected subgraph $G[V_{S}]$ of the molecular graph $G$. We call $V_{S}$ the source set. $G[V_{S}]$ simulates a valid molecular fragment with atom types $\mathbf{a}[V_{S}]=a_{1:n}$ and positions $\mathbf{X}[V_{S}]=\mathbf{x}_{1:n}$ in the autoregressive build-up at step $n=|V_{S}|$. The target set $V_{T}$ is the set of valid token continuations at step $n+1$. It is the set of boundary nodes $\partial V_{S}$, given $V_{S}$, and induces atom types $\mathbf{a}[V_{T}]$ and positions $\mathbf{X}[V_{T}]$ for model supervision.
The sampling of connected subgraphs is performed dynamically during training. First, we precompute the eccentricity $\epsilon(v)$ for each node $v\in V$ in the molecular graph. During runtime, a random seed node $v_{0}$ is selected from each molecule in the batch. Starting from this seed atom, we perform $m$ rounds of 1-hop neighborhood aggregation, where $m$ is a random number between 0 and $\epsilon(v_{0})$. This process generates connected subgraphs of varying sizes, ranging from a single node to the entire molecular graph.
To avoid biasing the sampling process toward globular subgraphs, we introduce stochasticity by randomly dropping a subset of the boundary nodes before aggregation, controlled by a dropout factor $\gamma$. To counteract the resulting bias toward smaller subgraphs, we introduce a scaling parameter $\beta$, which adjusts the number of growth iterations per graph. By balancing these two parameters, our sampling strategy generates connected subgraphs of diverse sizes and shapes, effectively simulating molecular growth during autoregressive generation. The pseudo-code for our sampling strategy is provided in Algorithm 1. In practice, this sampling strategy is efficiently implemented using batched operations and introduces only minimal overhead to the training process, which is dominated by the forward passes through the transformer.
@@
$m=\beta\cdot\text{RandomUniform}(0,\epsilon(v_{0}))$
for $i=0$ to m1m-1 do
  $\partial V_{S}\leftarrow(\bigcup_{i\in V_{S}}\mathcal{N}(i))\setminus V_{S}$
  $\partial V_{S}\leftarrow\{v\in\partial V^{\prime}\mid\text{RandomUniform}(0,1)>\gamma\}$
  $V_{S}\leftarrow V_{S}\cup\partial V_{S}$
@@
$V_{T}\leftarrow(\bigcup_{i\in V_{S}}\mathcal{N}(i))\setminus V_{S}$
@@
The computation of the source set representation $f_{\theta}:(\textbf{a}_{1:n},\textbf{X}_{1:n})\rightarrow\mathbf{z}_{n}$ at step $n$ is straightforward. Atom types $\textbf{a}_{1:n}$ are encoded using a linear embedding layer, while their positions $\mathbf{X}_{1:n}$ are encoded with Fourier features (tancik2020fourier). The summation of these encodings produces the initial node representations. For the transformer backbone $f_{\theta}$, we adopt the NanoGPT implementation (Karpathy2025nanoGPT) of GPT-2 (radford2019language), removing positional sequence embeddings and replacing causal attention with bidirectional attention. After the transformer updates, the node representations are aggregated by additive pooling (zaheer2017deep) to form a set-level representation $\mathbf{z}_{n}$, which characterizes the current source set at step $n$.
@@
where $p(a_{n+1}\mid\mathbf{z}_{n})$ represents the distribution over the next atom types, and $p(\mathbf{x}_{n+1}\mid a_{n+1},\mathbf{z}_{n})$ models the distribution over the next atom positions given the next atom type. The next atom type probability $p(a_{n+1}\mid\mathbf{z}_{n})$ is modeled as a categorical distribution over the vocabulary of atom types and is approximated with a linear layer followed by a softmax operation $g_{\theta}:\mathbf{z}_{n}\rightarrow\hat{a}_{n+1}$. For model supervision, we compute the mean over the one-hot-encoded atom types in the target set $\hat{a}_{n+1}=\text{mean}(\text{onehot}(\mathbf{a}[V_{T}]))$ and apply a cross-entropy loss.
The next atom position is modeled as a conditional distribution given the next atom type. We model this distribution via flow matching (Lipman2023FlowMatching; Albergo2023NormalizingFlows) (see Appendix Section A.1 for a more detailed description of the formalism). During training, we employ teacher forcing to decouple the atom type prediction from the task of positional recovery. For the set of possible next positions $\mathbf{X}_{1}=\mathbf{X}[V_{T}]$ corresponding to the target set $V_{T}$, we sample a set of random initial positions $\mathbf{X}_{0}\sim p_{0}$, where $p_{0}=\mathcal{N}(0,\sigma^{2}I)$. Following the flow matching convention, we construct interpolation paths between the positions at the random position at $t=0$ and the true position at $t=1$:

### v2

@@ Abstract
Transformer-based autoregressive models offer a promising alternative to diffusion- and flow-matching approaches for generating 3D molecular structures. However, standard transformer architectures require a sequential ordering of tokens, which is not uniquely defined for the atoms in a molecule. Prior work has addressed this by using canonical atom orderings, but these do not ensure permutation invariance of atoms, which is essential for tasks like prefix completion. We introduce NEAT, a Neighborhood-guided, Efficient, Autoregressive, Set Transformer that treats molecular graphs as sets of atoms and learns an order-agnostic distribution over admissible tokens at the graph boundary. NEAT achieves state-of-the-art performance in autoregressive 3D molecular generation whilst ensuring atom-level permutation invariance by design.
Machine Learning, ICML, 3D Molecular Generation, Set Transformer, Autoregressive Model, Flow Matching, Computer-Aided Drug Design
@@ 1 Introduction
Figure 1: Autoregressive generation of 3D molecules with NEAT. (1) A set transformer generates an abstract representation $\mathbf{z}_{n}$ of the molecule at state $n$, given the atom types $\mathbf{a}_{1:n}$ and the coordinates $\mathbf{X}_{1:n}$. (2) The next atom type $a_{n+1}$ is inferred from $\mathbf{z}_{n}$. (3) Conditioned on $a_{n+1}$ and $\mathbf{z}_{n}$, the next atom position $\mathbf{x}_{n+1}$ is modeled via a flow head, which concludes the molecule update at state $n+1$.
@@
Computational workflows for discovering bioactive molecules often rely on either virtual screening (Warr et al., 2022), or the direct de novo generation of molecules with desired properties (Peng et al., 2023). While the continuous expansion of chemical spaces increases the likelihood of identifying optimal designs through virtual screening, this growth also leads to increasingly higher computational costs (Bellmann et al., 2022). Moreover, despite their ultra-large scale, virtual chemical spaces still represent only a small fraction of the vast, theoretically possible molecular design space.
@@
Generative modeling offers a promising alternative to virtual screening by enabling the computationally cheaper generation of new molecules based on user-defined constraints. Molecular generation can be approached through the generation of (1D) SMILES strings (Kusner et al., 2017; Gómez-Bombarelli et al., 2018), (2D) molecular graphs (Simonovsky and Komodakis, 2018; Cao and Kipf, 2022), (3D) atomic point clouds (Gebauer et al., 2019; Satorras et al., 2021; Hoogeboom et al., 2022; Luo and Ji, 2022; Xu et al., 2023; Daigavane et al., 2024; Zhang et al., 2025; Cheng et al., 2025), or a combination thereof (Hua et al., 2023; Peng et al., 2023; Vignac et al., 2023). Molecules are inherently three-dimensional, with many properties determined by their conformations. While modeling 3D structures is challenging, the additional spatial information is useful for many downstream applications. In this work, we focus on unconditional 3D generation, which is a necessary prerequisite to condition the generation process towards compounds with specific biological functions later on (Peng et al., 2022; Schneuing et al., 2024).
@@
State-of-the-art 3D molecular generation models rely on diffusion and flow matching techniques (Hoogeboom et al., 2022; Xu et al., 2023; Song et al., 2023, 2024; Zhang et al., 2025; Feng et al., 2025; Dunn and Koes, 2025). A key limitation of these methods is the need to compute the joint atom distribution at each diffusion step through computationally intensive, quadratically scaling message-passing updates in the encoder model to capture atom interdependencies. Additionally, diffusion and flow matching molecular generators require the number of atoms to be predefined. While the use of dummy atoms offers a partial workaround (Adams et al., 2025; Dunn and Koes, 2025), these approaches remain constrained by the upper bound on the number of nodes sampled from the noise distribution.
@@
Autoregressive models offer a flexible and efficient alternative. They support variable atom counts and require only one encoder call per generated token (Li et al., 2024). For 3D autoregressive molecular generation, there currently exist three lines of work: (1) using graph neural networks (GNNs) in combination with focus atoms (Gebauer et al., 2019; Luo and Ji, 2022; Daigavane et al., 2024); (2) using standard large language models (LLMs) for generating XYZ, CIF or PDB files (Flam-Shepherd and Aspuru-Guzik, 2023) or SE(3)-invariant 1D tokens (Li et al., 2025b; Fu et al., 2025; Lu et al., 2025); (3) combining a causal transformer with a diffusion loss to generate continuous tokens (atomic coordinates) (Cheng et al., 2025; Li et al., 2025a). Although current GNN-based methods ensure permutation invariance with respect to atom ordering, they generate molecules with low chemical validity. Conversely, sequence-based methods perform well in terms of validity but are not permutation-invariant with respect to atom ordering.
@@
In this work, we present the first autoregressive transformer model for 3D molecular generation that respects the atom permutation-invariance of molecular graphs. Working without canonical atom orderings introduces two challenges. First, it precludes the use of sequence embeddings, which are an essential component of standard autoregressive models. Second, it changes the autoregressive training set-up from supervision with a single token per step to supervision with a set of valid tokens: without a canonical atom order, all atoms in the 1-hop neighborhood of the current graph are valid continuations.
@@
Our solution to these challenges is simple yet effective. Given a partial molecular graph, a permutation-invariant set transformer (Lee et al., 2019) is trained to produce a token distribution that matches the boundary nodes of the subgraph (i.e., all 1-hop neighbors of existing nodes). This training signal, which we refer to as neighborhood guidance, yields a probabilistic next-token distribution without relying on ad hoc canonicalization schemes. We combine it with a flow matching loss to model continuous atomic coordinates.
Our model, NEAT (Neighborhood-guided, Efficient Autoregressive set Transformer), matches state-of-the-art autoregressive 3D molecular generation models on the QM9 and GEOM-Drugs benchmarks while preserving permutation invariance with respect to atom ordering. The permutation invariance enables robust prefix completion without any modifications to the model, yielding unprecedented generation quality.
@@ 2 Related works
$p(x)=p(x_{1})\prod_{n=1}^{N-1}p(x_{n+1}\mid x_{1:n}),$ (1)
where $x_{1:n}=(x_{1},\dots,x_{n})$ represents the tokens up to step $n$. Sampling is performed by initializing with $x_{1}$ and sequentially generating tokens until $x_{N}$. This is straightforward for sequentially ordered data such as text or images. In the graph domain, the generation order is less obvious, and has been an active focus of research. One way to model graphs as sequences is through canonicalization. However, predefined canonical orderings impose limitations, as the optimal generation order is often context-dependent. To alleviate strict assumptions, You et al. (2018) proposed a breadth-first search node ordering for autoregressive graph generation, while Wang et al. (2025b) introduced order policies to dynamically assign generation orders based on the current state. Although these approaches reduce reliance on fixed orderings, they do not fully eliminate order dependence. 3D molecules are geometric graphs whose structure is determined by connectivity and geometry. The generation of valid positions poses additional constraints on the graph generation process and recent work on this specific task will be discussed in the following.
@@
G-SchNet (Gebauer et al., 2019) is an $E(3)$-invariant GNN model that introduces the concept of focus atoms for generation order. In each generation step, the model selects a focus atom as the center of a 3D grid, chooses the next atom type, and infers its position probability from radial distances to previously placed atoms. G-SphereNet (Luo and Ji, 2022) adopts this approach but places atoms relative to the focus atom by generating atom distances, bond angles, and torsion angles with a normalizing flow. The $E(3)$-equivariant Symphony model (Daigavane et al., 2024) parameterizes the 3D probability around the focus atom using $E(3)$-equivariant features from spherical harmonic projections rather than rotationally invariant features. While all these models respect the permutation-invariance of atoms, they produce molecules with comparably low chemical validity.
Transformer models with discrete tokenization
Transformers can be used for 3D molecule generation when combined with a suitable text format. Flam-Shepherd and Aspuru-Guzik (2023) trained an LLM for 3D molecular generation directly on XYZ, CIF, or PDB files. An active line of research are tokenization schemata to allow LLMs more fine-grained control in the generation process. Geo2Seq (Li et al., 2025b) converts molecular geometries into $SE(3)$-invariant 1D discrete sequences for unconditional modeling, while Frag2Seq (Fu et al., 2025) uses a related technique for structure-based modeling. Uni-3DAR (Lu et al., 2025) employs a voxel-based discretization scheme with an octree-based compression technique, yielding sparse sequences.
@@
One of the challenges in generating 3D molecular structures with transformers is to generate continuous-valued tokens like 3D atomic coordinates without discrete tokenization strategies. While standard transformer architectures were initially developed for categorical data, they can be adapted for the generation of continuous values. Li et al. (2024) directly modeled the next-token distribution $p(x_{n+1}\mid x_{1:n})$ using a diffusion-based approach, factorized as:
$p(x_{n+1}\mid x_{1:n})=p(x_{n+1}\mid\mathbf{z}_{n})\cdot p(\mathbf{z}_{n}\mid x_{1:n}),$ (2)
where $\mathbf{z}_{n}\in\mathbb{R}^{k}$ represents an abstract latent variable for the next token. This factorization enables to use a transformer to model $p(\mathbf{z}_{n}\mid x_{1:n})$, capturing interdependencies among tokens, followed by a lightweight diffusion-based multilayer perceptron (MLP) to map $\mathbf{z}_{n}$ to the continuous token space. Their approach enabled high-quality autoregressive image generation with improved inference speed.
Cheng et al. (2025) applied this idea to the generation of atomic positions, enabling 3D molecular generation without spatial discretization. They employed a causal transformer with sequential embeddings derived from the atom order in XYZ coordinate files to predict discrete atom types, and a diffusion model, conditioned on these predictions, to predict the corresponding continuous 3D atom positions. However, the study also shows that reliance on a causal transformer introduces limitations. The use of sequence information based on atom order in XYZ files makes the model susceptible to overfitting and degrades performance when the atom order is perturbed. Li et al. (2025a) adopted a comparable framework, adding inertial-frame alignment for coordinates and canonical atom ordering, but their method relocates the problem of token-order sensitivity rather than eliminating it.
We propose that a set transformer (Lee et al., 2019) with bidirectional attention is better suited for autoregressive 3D molecular generation, as it inherently respects the permutation invariance of atom sets by disregarding the generation order of atoms at state $n$. This raises the challenge of compensating for the absence of sequential order information, which is the key focus of our work.
@@ 3 Methods
Our architecture consists of three main components: (1) a data-augmentation module that provides neighborhood-guidance during training, (2) a set transformer trunk that produces permutation-invariant encodings of atom sets, and (3) a flow matching head that generates continuous atomic coordinates. During training, we split molecules into a source set, representing the state of the molecular graph at step $n$, and a target set, representing the set of possible token continuations at step $n+1$. The source set is processed by the set transformer to obtain a learned representation of the current state. This source set representation is then passed through a simple linear layer followed by a softmax to yield a categorical distribution over the next atoms’ types. The continuous next atoms’ positions are modeled by the flow matching head, conditioned on the source set representation and the next atom type. During training, we use teacher forcing to decouple the prediction of atom types from the prediction of atom positions. Because the expensive source set representation is computed only once per generated token, our model is computationally more efficient than non-autoregressive frameworks that rely entirely on diffusion or flow matching. Training and inference procedures are summarized in Figure 2.
@@
Figure 2: (1) Overview of the proposed training workflow. The model takes molecular graphs as input, where nodes encode atom types and positions, and edges represent chemical bonds. For each training example, we randomly sample a connected subgraph. The selected nodes form the source set, and their neighboring boundary nodes define the target set. The source set is encoded using a set transformer, and its representation is used to predict the distribution over the next atom type. For each target-set node, we sample a position from a normal distribution and construct linear interpolation paths between the sampled and true positions. These interpolation paths provide supervision for learning the vector field via flow matching. (2) Overview of the proposed model inference workflow.
@@
In the build-up phase, the model iteratively predicts atom type $a_{n+1}$ and position $\mathbf{x}_{n+1}$, given the current set of atoms and positions at step $n$. We will use the following compact notation $\mathbf{a}_{1:n}=(a_{1},\dots,a_{n})$ to refer to the generated atom types and positions $\mathbf{X}_{1:n}=(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ at step $n$. The autoregressive build-up stops when the model predicts a stop token. Edges are then inferred from the $N$ generated atoms via a predefined set of rules like $E=\mathcal{R}(\mathbf{a},\mathbf{X})$, which completes inference of the molecular graph $G$. While edges are inferred from atom types and positions after the build-up phase, the edge information can still be used for teacher guidance in the training phase. In the following, we describe how we use this information for model training.
@@
We propose a sampling strategy that mimics molecular growth during the autoregressive build-up phase. Given a molecular graph $G$ in the training set, we sample a node-induced, connected subgraph $G[V_{S}]$ of $G$. We call $V_{S}$ the source set. $G[V_{S}]$ simulates a molecular fragment with atom types $\mathbf{a}[V_{S}]=a_{1:n}$ and positions $\mathbf{X}[V_{S}]=\mathbf{x}_{1:n}$ in the autoregressive build-up at step $n=|V_{S}|$. The target set $V_{T}$ is the set of valid token continuations at step $n+1$. It is the set of boundary nodes $\partial V_{S}$, given $V_{S}$, and induces atom types $\mathbf{a}[V_{T}]$ and positions $\mathbf{X}[V_{T}]$ for model supervision.
The sampling of connected subgraphs is performed dynamically during training. First, we compute the eccentricity $\epsilon(v)$ (the maximum shortest-path distance from $v$ to all other nodes in $G$) for each node $v\in V$. During runtime, a random seed node $v_{0}$ is selected from each molecular graph in the batch. Starting from this seed node, we perform $m$ rounds of 1-hop neighborhood aggregation, where $m$ is a random integer between $1$ and $\epsilon(v_{0})$. To avoid biasing the sampling process toward globular subgraphs, we introduce stochasticity by randomly dropping a subset of the boundary nodes in each aggregation round, controlled by a dropout factor $\gamma$. To counteract the resulting bias toward smaller subgraphs, we introduce a scaling parameter $\beta$, which adjusts the number of aggregation rounds per graph. By balancing these two parameters, our sampling strategy generates connected subgraphs of varying sizes and shapes, from a single node to the full molecular graph, thereby simulating molecular growth in an autoregressive manner.
The pseudo-code for our sampling strategy is provided in Algorithm 1. Additional information on the impact of $\gamma$ and $\beta$ on the sampling strategy and the values used in this work are available in Appendix A.3. In practice, this sampling strategy is efficiently implemented using batched operations and introduces only minimal overhead to the training process, which is dominated by the forward passes through the transformer.
@@
$m=\lfloor{\beta\cdot\epsilon(v_{0})\cdot u_{0}}\rfloor+1,\;u_{0}\sim\text{Uniform}([0,1))$
for $i=1$ to $m$ do
  $\partial V_{S}\leftarrow(\bigcup_{j\in V_{S}}\mathcal{N}(j))\setminus V_{S}$
  $\partial V_{S}^{\prime}\leftarrow\{v\in\partial V_{S}\mid u_{i}>\gamma,u_{i}\sim\text{Uniform}(0,1)\}$
  $V_{S}\leftarrow V_{S}\cup\partial V_{S}^{\prime}$
@@
$V_{T}\leftarrow(\bigcup_{j\in V_{S}}\mathcal{N}(j))\setminus V_{S}$
@@
The computation of the source set representation $f_{\theta}:(\mathbf{a}_{1:n},\mathbf{X}_{1:n})\rightarrow\mathbf{z}_{n}$ at step $n$ is straightforward. Atom types $\mathbf{a}_{1:n}$ are encoded using a linear embedding layer, while their positions $\mathbf{X}_{1:n}$ are encoded with Fourier features (Tancik et al., 2020). The summation of these encodings produces the initial node representations. Note that the source set atom positions are not re-centered prior to the computation of the source set representation. For the transformer backbone $f_{\theta}$, we adopt the NanoGPT implementation (Karpathy, 2025) of GPT-2 (Radford et al., 2019), removing positional sequence embeddings and replacing causal attention with bidirectional attention. After the transformer updates, the node representations are aggregated by additive pooling (Zaheer et al., 2017) to form a set-level representation $\mathbf{z}_{n}$, which characterizes the current source set at step $n$.
@@
where $p(a_{n+1}\mid\mathbf{z}_{n})$ represents the distribution over the next atom types, and $p(\mathbf{x}_{n+1}\mid a_{n+1},\mathbf{z}_{n})$ models the distribution over the next atom positions given the next atom type. The next atom type probability $p(a_{n+1}\mid\mathbf{z}_{n})$ is modeled as a categorical distribution over the vocabulary of atom types and is approximated with a linear layer followed by a softmax operation $g_{\theta}:\mathbf{z}_{n}\rightarrow\hat{a}_{n+1}$. For model supervision, we compute the cross-entropy loss between $\hat{a}_{n+1}$ and the observed distribution over target atom types, which is obtained by computing the mean over the one-hot-encoded atom types in the target set $a_{n+1}=\text{mean}(\text{onehot}(\mathbf{a}[V_{T}]))$.
The next atom position is modeled as a conditional distribution given the next atom type. We model this distribution via conditional flow matching (CFM, see Appendix A.1 for a more detailed description of the formalism) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023). During training, we employ teacher forcing to decouple the atom type prediction from the task of positional recovery. For the set of possible next positions $\mathbf{X}_{1}=\mathbf{X}[V_{T}]$ corresponding to the target set $V_{T}$, we sample a set of random initial positions $\mathbf{X}_{0}\sim p_{0}$, where $p_{0}=\mathcal{N}(0,\sigma^{2}I)$. We construct interpolation paths between the positions at the random position at $t=0$ and the true position at $t=1$: